OpenStudy (anonymous):

What is the equation of the line that passes through (4, -1) and (-2, 3)?

OpenStudy (anonymous):

First of all, remember what the equation of a line is: y = mx+b

OpenStudy (anonymous):

right

OpenStudy (anonymous):

m=slope b=y intercept

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

rise over run = 2/2?

OpenStudy (anonymous):

For lines like these, the slope is always defined as "the change in y over the change in x" or, in equation form: \[m=y2 - y1 / x2 - x1 \] Once you plug in the information, the slope you should get is -2/3

OpenStudy (zzr0ck3r):

(3-(-1))/(-2-4)

OpenStudy (anonymous):

Now let's find the yintercept!

OpenStudy (anonymous):

ok!

OpenStudy (anonymous):

The b is what we want. We can plug anything we want in for x and y here, but we want the equation for the line that specfically passes through the two points (4,-1) and (-2,3).

OpenStudy (anonymous):

Plug in for x and y from one of our (x,y) points that we know the line passes through.

OpenStudy (anonymous):

This will allow us to solve for b for the particular line that passes through the two points you gave! You can use either (x,y) point you want..the answer will be the same: (4, -1). y=mx+b or -1=-2/3 *4+b or solving for b: b=-1-(2/3)(4) b=5/3

OpenStudy (anonymous):

OR

OpenStudy (anonymous):

(-2,3). y=mx+b or 3=-2/3*2+b or solving for b: b=3 -(-2/3)(-2) b=5/3 Same answer!

OpenStudy (anonymous):

This is what will complete our problem.

OpenStudy (anonymous):

The final answer will be: y=-2/3x+5/3

OpenStudy (anonymous):

Ok im sorta understanding it.

OpenStudy (anonymous):

=) Hope this helped!

OpenStudy (anonymous):

Good. Just keep practicing. You'll get it!

OpenStudy (anonymous):

Wait how would that look in general form because the awnser to my problem is compley different

OpenStudy (anonymous):

Answers*

OpenStudy (anonymous):

The "General Form" of the equation of a straight line is: Ax + By + C = 0

OpenStudy (anonymous):

Is that what they look like on your answers?

OpenStudy (anonymous):

Step 1: Calculate the slope (m) from the coordinates: (y2 - y1) / (x2 - x1) and reduce the resulting fraction to the simplest form. Step 2: From the slope, calculate variables A and B with the equation: Slope = - A / B Step 3: Calculate the variable C by applying one of the coordinates to the equation: Ax + By = -C Result: Now you have calculated all three variables (A, B and C) for the General Form Linear Formula. For your example: -2x - 3y +5 =0

OpenStudy (anonymous):

Good luck!

OpenStudy (anonymous):

=)